Determine the distance R to B for the aunch conditions shown. Ans: R 2386.37m A-150 m/s 150 m
Answer #1

Let us consider the upwards direction as positive and the downwards direction as negative.

Gravitational acceleration = g = -9.81 m/s2

Initial velocity of the projectile = V = 150 m/s

Angle of launch = \theta

From the figure,

Sin\theta = 3/5 = 0.6

Cos\theta = 4/5 = 0.8

Initial horizontal velocity of the projectile = Vx = VCos\theta = (150)(0.8) = 120 m/s

Initial vertical velocity of the projectile = Vy = VSin\theta = (150)(0.6) = 90 m/s

Initial height of the point from where it is launched = H = 150 m

Time taken by the projectile to reach the ground = T

When the projectile lands on the ground the displacement of the projectile in the vertical direction is downwards therefore it is negative.

-H = VyT + gT2/2

-150 = (90)T + (-9.81)T2/2

4.905T2 - 90T - 150 = 0

-b + V6 - 4ac T=- 2a

(-90)+ (-9012 - 44.90511-150) 214.905)

T = 90 + 105.0856 9.81

T = 19.8864 sec or -1.538 sec

Time cannot be negative.

T = 19.8864 sec

There is no horizontal force on the projectile therefore the horizontal velocity of the projectile remains constant.

Horizontal displacement of the projectile = R

R = VxT

R = (120)(19.8864)

R = 2386.37 m